I do not understand how to solve equations. eample
-2x+8=104+6x
People have tried to explain to me and when I think I have it figured out, it's wrong. I need a simple way of solving these types of equations.. "HELP"

Answer by nyc_function(2741)

Just combine like terms. In other words, put the x terms together and the constants together.
Keep in mind that the sign in front of each term changes when crossing the equal sign.
-2x + 8 = 104 + 6x

-2x - 6x = 104 - 8
-8x = 96
x = 96/-8
x = -12